I can't seem to get this one at all...
-11+10(p+10)=4-5(2p+11)

Answer by stanbon(75887)

-11+10(p+10)=4-5(2p+11)
-----
-11 + 10p + 100 = 4 -10p - 55
------
20p + 89 = -51
----
20p = -140
----
p = -7
===============
Cheers,
Stan H.
===============